### v2

# Smooth Tests for Normality in ANOVA

Abstract
The normality assumption for random errors is fundamental in the analysis of variance (ANOVA) models, yet it is rarely formally tested in practice. In this paper, we propose Neyman’s smooth tests for assessing the normality assumption across various types of ANOVA models. The proposed test statistics are constructed based on the Gaussian probability integral transformation of ANOVA residuals. Under the null hypothesis of normality, the test statistics are asymptotically Chi-square distributed, with degrees of freedom determined by the dimension of the smooth model (the number of orthonormal functions). A data-driven selection of the model dimension using a modified Schwarz’s criterion is also discussed. Simulation studies demonstrate the effectiveness of our proposed method.
Keywords: ANOVA; Estimation effect; Normality; Schwarz’s selection rule; Smooth test.
1 Introduction
Analysis of Variance (ANOVA) is an important and widely used method in exploratory and confirmatory data analysis (Gelman,, 2005), particularly for comparing group means and assessing the significance of factors in experimental designs. The theory of ANOVA models has been well established and developed in the literature, see, for example, Scheffé, (1959), Miller Jr, (1997), Wickens and Keppel, (2004), Dean et al., (2017), Hirotsu, (2017), Montgomery, (2017). A standard assumption in ANOVA models is that the random errors are normally distributed. Nonnormality may make the normal-based inference invalid. In terms of estimation, departures from normality generally undermine the validity of the normal-theory variance formulae for point estimates of variance components, because these formulae rely on theoretical results that represent the mean squares of random effects as constant multiples of Chi-square variables (Scheffé,, 1959). From the view of testing, the $F$-test employed in ANOVA is sensitive to nonnormality if the group sizes are unequal (Ali and Sharma,, 1996). The effects of nonnormality on size and power of the $F$-test are also extensively investigated in details (Pearson,, 1931; Gayen,, 1950; David and Johnson, 1951a, ; David and Johnson, 1951b, ; Srivastava,, 1959; Atiqullah,, 1962; Tiku,, 1964, 1971; Donaldson,, 1968; Janic-Wróblewska and Ledwina,, 2000). Therefore, it is crucial to evaluate the normality of ANOVA models.
While a broad range of normality tests have been proposed for observed data and linear models, comparatively little effort has been devoted to systematically checking the normality assumption in ANOVA. A direct approach to normality diagnosis is to plot a normal probability plot of ANOVA residuals (Dean et al.,, 2017; Hirotsu,, 2017), which is intuitive but lacks a theoretical guarantee. To our knowledge, formal hypothesis testing procedures for assessing normality in ANOVA appear to be relatively scarce in the existing literature. Existing methods are merely a simple application of classical normality tests to ANOVA residuals, see, for instance, Bonett and Woodward, (1990), Bonett and Seier, (2002) and Hwang and Wei, (2006). Although applicable in practice, these methods do not precisely provide a rigorous assessment of normality within the ANOVA framework, nor do they account for the specific structure and constraints inherent in ANOVA models.
In this article, following the spirit of Neyman, (1937), we propose a unified smooth test framework for assessing the normality assumption in various types of ANOVA models. Neyman’s smooth test has gained popularity across various scientific disciplines due to its well-known advantages in both theoretical and practical aspects. It has proven to be powerful in detecting against a broad class of alternatives; see, e.g., Bera and Ghosh, (2002), Thas, (2010), and Lehmann and Romano, (2022). Specifically, we first re-express the problem of testing for normality as testing for uniformity via the Gaussian probability integral transformation (PIT), and then construct the test statistics through the Gaussian PIT of ANOVA residuals. Under mild conditions for each model, the proposed test statistics are asymptotically Chi-square distributed under the null. Furthermore, we analyze the power properties of the tests under fixed and local alternatives. A modified Schwarz’s selection rule is proposed to determine the order of tests. Monte Carlo experiments demonstrate the good performance of our methodology in finite samples.
The advantages and contributions of our proposed tests are as follows. First, the proposed asymptotic Chi-square tests are simple to implement, in the sense that the critical values can be obtained directly from the limiting Chi-square distributions without bootstrapping. Second, we systematically investigate the properties of test statistics for three types of one-way fixed effects models with potentially diverging numbers of groups in a unified asymptotic framework. The parameter estimation effects in each model are well interpreted through our analysis procedure. The heterogeneity in group means and group variances does contribute to the specific form of test statistics and to the corresponding regularity conditions, but does not alter the convergence rate, which is only related to the total sample sizes under both the null and the alternatives. Lastly, we clarify the motivation and theoretical properties of data-driven smooth tests under ANOVA settings, and provide a revised limiting null distribution of the data-driven test statistics in finite samples.
The rest of this paper is organized as follows. Section 2 presents the general testing framework. Section 3 establishes the theoretical results for smooth tests in three types of ANOVA models. In Section 4, a data-driven approach is proposed to determine the order of tests. Simulation results in Section 5 demonstrate good finite-sample performance of our proposed methodology. Some extensions are discussed in Section 6. Section 7 concludes. The proof of theoretical results and some practical issues are gathered in the supplementary material.
2 The testing framework
Our problem of interest is whether the random error $\varepsilon=\sigma e$ in ANOVA models is normally distributed with mean zero and variance $\sigma^{2}$ (denoted as $\mathcal{N}(0,\sigma^{2})$) for some $\sigma>0$. We aim to test for the null hypothesis $H_{0}:e\sim\mathcal{N}(0,1)$ against the alternatives of nonnormality. To motivate our methodology, consider the transformed random variable $Z=\Phi(e)$, with $\Phi(\cdot)$ denoting the CDF of the standard normal distribution. Under the null hypothesis of normality, the CDF of $Z$ is given by
$G(z)=\mathrm{P}(Z\leq z)=\mathrm{P}\left(e\leq\Phi^{-1}(z)\right)=\Phi(\Phi^{-1}(z))=z.$
Or, equivalently, under $H_{0}$, the probability density function (PDF) of $Z$ is $g(z)\equiv 1$ for $z\in[0,1]$. That is, $Z\sim\mathcal{U}[0,1]$ under $H_{0}$, where $\mathcal{U}[0,1]$ marks the uniform distribution on the interval $[0,1]$. Under the alternatives, the PDF $g(z)$ differs from 1. The distinct behavior of $Z$ under $H_{0}$ and $H_{1}$ provides the basis for Neyman’s smooth test. Neyman, (1937) considered the following smooth alternative to the uniform density:
$h\left(z\right)=c\left(\bm{\theta}_{K}\right)\exp\left[\sum_{k=1}^{K}\theta_{k}\pi_{k}\left(z\right)\right],\quad 0\leq z\leq 1,$ (2.1)
where $c(\bm{\theta}_{K})$ is a constant of integration depending on $\bm{\theta}_{K}=(\theta_{1},\theta_{2},...,\theta_{K})^{\top}$, and $\{\pi_{k}\left(z\right)\}_{k=0}^{\infty}$ is an orthonormal system in $L_{2}[0,1]$ with $\pi_{0}(z)=1$ and
[…]
The null hypothesis that $Z\sim\mathcal{U}[0,1]$ can then be tested by testing $\theta_{1}=\theta_{2}=\ldots=\theta_{K}=0$ in (2.1).
If we had independent and identically distributed (i.i.d.) observations $\{\varepsilon_{i}\}_{i=1}^{n}$ and $\sigma$ were known, the smooth test statistic for testing $H_{0}$ has the following quadratic form:
$\Psi_{K}^{2}=\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}(Z_{i})\right)^{2},$ (2.3)
where $K$ is some fixed positive integer and $Z_{i}=\Phi(e_{i})=\Phi(\varepsilon_{i}/\sigma)$. Under $H_{0}$, $n\Psi_{K}^{2}$ converges asymptotically to a Chi-square distribution with $K$ degrees of freedom (denoted as $\chi_{K}^{2}$), and the test has the local optimal properties of Rao’s score test.
However, it is always the case that $\{\varepsilon_{i}\}_{i=1}^{n}$ is unobserved and $\sigma$ is unknown. We may use some approximation $\{\hat{\varepsilon}_{i}\}_{i=1}^{n}$ for $\{\varepsilon_{i}\}_{i=1}^{n}$ and some estimator $\hat{\sigma}$ for $\sigma$ (e.g., the residuals and the sample standard deviation, respectively). Then, consider the following approximation for $Z_{i}$:
$\hat{Z}_{i}=\Phi(\hat{e}_{i})=\Phi\left(\frac{\hat{\varepsilon}_{i}}{\hat{\sigma}}\right),\quad i=1,\ldots,n.$
With $\{\hat{Z}_{i}\}_{i=1}^{n}$ replacing $\{Z_{i}\}_{i=1}^{n}$, we can construct a feasible smooth test statistic of the following form:
$\sum_{k=1}^{K}\left(\frac{1}{n}\sum_{i=1}^{n}\pi_{k}\left(\hat{Z}_{i}\right)\right)^{2}.$
Due to the replacement of $Z_{i}$ by $\hat{Z}_{i}$, the sequence $\{\hat{Z}_{i}\}_{i=1}^{n}$ may not be i.i.d. any longer. Therefore, it is crucial to study the asymptotic properties of ${n}^{-1}\sum_{i=1}^{n}\pi_{k}(\hat{Z}_{i})$. Compared with $\Psi_{K}^{2}$ in (2.3), $\hat{Z}_{i}$ has non-negligible estimation effects, as demonstrated in the theoretical results below. Consequently, the presence of estimation effects invalidates the above form of the test statistic and requires a different normalization matrix to restore the $\chi_{K}^{2}$ limiting null distribution. This motivates us to study the estimation effects carefully and then derive the correct quadratic-form test statistics.
It is also important to address the issue of sample sizes in ANOVA models. Throughout this section, to highlight the essence of smooth tests, we do not directly consider the ANOVA models but first examine the testing for a single random sequence of sample size $n$. This simplified setting captures the core ideas of the methodology and facilitates the derivation of key asymptotic properties. In the complete ANOVA framework, additional complications arise due to multiple sample sizes and the indexing structure, which require a more elaborate treatment. Besides, from a practical perspective, our method is expected to accommodate scenarios in which the number of groups grows with the total sample size. In subsequent analysis, we will systematically tackle these issues for specific ANOVA models.
3 Testing for normality in one-way fixed effects models
In this section, we establish smooth tests for normality in three types of one-way fixed effects models where either the group means or the group variances are equal. We examine the testing procedure and the theoretical properties separately for each case. Let $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ denote the observations of interest coming from $J$ groups, with $N_{j}$ being the sample sizes in group $j$ for $1\leq j\leq J$. Define $N=\sum_{j=1}^{J}N_{j}$ as the total sample size. In our theoretical framework, the number of groups $J$ is allowed to diverge as $N$ increases.
3.1 Same mean and same variance
We first investigate the ANOVA models with equal group means and equal group variances. Consider the following model:
$Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.1)
where $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. random errors with mean zero and variance one, $\mu$ is the common group mean, and $\sigma$ is the standard deviation of random errors. We also define $Z_{ij}=\Phi(e_{ij})$.
We estimate $\mu$ and $\sigma^{2}$ in (3.1) by the sample mean $\hat{\mu}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}Y_{ij}$ and the sample variance $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}\right)^{2}$, respectively. Define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu})/\hat{\sigma}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To derive the test statistic and its asymptotic distribution, we first analyze the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$. The following two assumptions are imposed for our asymptotic analysis.
Assumption 3.1.
$\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d. with continuous CDF $F(x)$, PDF $f(x)$, mean zero, variance one, and finite fourth moments.
Assumption 3.2.
For $k=1,\ldots,K$, $\pi_{k}(\cdot)$ are twice differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
The two mild conditions are in line with similar assumptions adopted in the literature of Neyman’s smooth tests, including Bera et al., (2013) and Song and Xiao, (2022). Introduce two constants:
$c_{1k}=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z,\quad c_{2k}=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z,\quad k=1,\ldots,K.$
Theorem 1.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $N\to\infty$.
Theorem 1 states that $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ is asymptotically equivalent to the sum of three terms: $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$, and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$. The latter two represent the estimation effects due to estimating $\mu$ and $\sigma^{2}$, respectively, and they contribute to the limiting distribution of the feasible smooth test statistic based on $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$, $k=1,\ldots,K$.
With the assistance of Theorem 1, by the central limit theorem (CLT), we can obtain the asymptotic normality of the $K$-dimensional vector
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$ (3.2)
$\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Sigma}_{K}\right),$
where $\bm{\Sigma}_{K}=\left(\sigma_{kl}\right)_{K\times K}$ is given by
$\sigma_{kl}$ $=\mathrm{E}\left[\left(\pi_{k}(Z)-c_{1k}e-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right)\left(\pi_{l}(Z)-c_{1l}e-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right)\right]$
$=\delta_{kl}-c_{1k}c_{1l}-\frac{1}{2}c_{2k}c_{2l}.$
Based on this fact, our proposed feasible smooth test statistic is defined as
$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\bm{\Sigma}_{K}^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.3)
The following corollary establishes the limiting null distribution of $\hat{\Psi}_{K}^{2}$ defined in (3.3).
Corollary 1.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2},$
as $N\to\infty$.
Corollary 1 demonstrates an asymptotic $\chi^{2}$ test for the normality of model (3.1) based on the statistic $N\hat{\Psi}_{K}^{2}$. Given the asymptotic significance level of $\alpha$, we reject the null hypothesis if
$N\hat{\Psi}_{K}^{2}>\chi_{K,1-\alpha}^{2},$
where $\chi_{K,1-\alpha}^{2}$ denotes the $(1-\alpha)$-th quantile of the $\chi_{K}^{2}$ distribution.
Now we turn to the asymptotic behavior of $\hat{\Psi}_{K}^{2}$ under the fixed alternatives and a Pitman-type sequence of local alternatives. For the fixed alternatives, we consider the following form:
[…]
Under $H_{1}$ given by (3.4), the random error $e$ is no longer normally distributed, which leads to different asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ from those under $H_{0}$. To formally present the corresponding theoretical results under $H_{1}$, we need some notations. Let
$d_{1k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)f(x)\mathrm{d}x,$
$d_{2k}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{-\infty}^{\infty}\dot{\pi}_{k}\left(\Phi\left(x\right)\right)\phi\left(x\right)xf(x)\mathrm{d}x,$
$d_{3k}=\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)xf(x)\mathrm{d}x,$
and
$d_{4k}=\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}\left(\Phi\left(x\right)\right)x^{2}f(x)\mathrm{d}x,$
for $k=1,\ldots,K$. Note that under $H_{0}$, $d_{1k}=d_{3k}=c_{1k}$ and $d_{2k}=d_{4k}=c_{2k}$ due to $f(x)=\phi(x)$. The following theorem demonstrates the asymptotic properties under $H_{1}$.
Theorem 2.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.5)
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $N\to\infty$. Furthermore,
$\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
and
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right),$ (3.6)
where $\bm{a}_{K}\equiv(a_{1},\ldots,a_{K})^{\top}=(\mathrm{E}[\pi_{1}(Z)],\ldots,\mathrm{E}[\pi_{K}(Z)])^{\top}$ and $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ is given by
$\xi_{kl}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
From Theorem 2, one can see that under $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
$a_{k}\equiv\mathrm{E}[\pi_{k}(Z)]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))f(x)\mathrm{d}x=\int_{0}^{1}\pi_{k}(z)\frac{f(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z\neq 0,$
the smooth test statistic $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, and the asymptotic power of the test is 1. Specifically, the power function is given by
$1-\Phi\left(\left(4N\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)^{-1/2}\chi_{K,1-\alpha}^{2}-\left(N\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}/4\right)^{1/2}\right).$
For further power analysis, we consider a Pitman-type sequence of local alternatives converging to the null hypothesis at a suitable rate, which is specified as follows:
$H_{1L}:F(x)=(1-\delta_{N})\Phi(x)+\delta_{N}Q(x),$ (3.7)
where $Q\left(x\right)$ (which admits PDF $q\left(x\right)$) represents some distribution function that is different from $\Phi(x)$, and $\delta_{N}\to 0$ as $N\to\infty$. Denote
$\Delta_{k}=\int_{0}^{1}\pi_{k}(z)\frac{q(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z,\quad k=1,\ldots,K.$
The following theorem presents the theoretical properties of $N\hat{\Psi}_{K}^{2}$ under the local alternatives. Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the local power of the test.
Theorem 3.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.8)
$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $N\to\infty$. Furthermore,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
where $\bm{\Delta}_{K}=(\Delta_{1},\ldots,\Delta_{K})^{\top}$, and $\chi^{2}_{K}\left(\tau\right)$ denotes the noncentral $\chi^{2}$ distribution with $K$ degrees of freedom and nonnegative noncentrality parameter $\tau$.
It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to have a nontrivial power against the Pitman sequence of local alternatives. It would be impossible to distinguish the alternatives from the null if $\delta_{N}=o(N^{-1/2})$. Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ has nontrivial asymptotic power against the local alternatives.
Remark 1.
We should emphasize that the proposed smooth test is not a “consistent” test in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives. Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power go to 1. Otherwise, if $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathrm{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\hat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$. For example, if $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ such that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and we take the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$, then for $Z=\Phi(e)$, we have
$\mathrm{E}[Z]$ $=\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}\Phi(z)\mathrm{d}z$
$=\frac{1}{2\sqrt{3}}z\Phi(z)\Big{|}_{-\sqrt{3}}^{\sqrt{3}}-\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}z\mathrm{d}\Phi(z)$
$=\frac{1}{2}\left(\Phi\left(\sqrt{3}\right)+\Phi\left(-\sqrt{3}\right)\right)$
$=\frac{1}{2},$
which leads to $\mathrm{E}[\pi_{1}(Z)]=0$. In fact, smooth tests are neither directional nor omnibus; that is, they maintain decent power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for details.
3.2 Different means and same variance
Next we allow the group means $\mu$ in model (3.1) to vary across the group index $j$, in the sense that
$Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.9)
where $\{\mu_{j}\}_{j=1}^{J}$ represents the potentially different means of each group. Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be estimators of $\mu_{j},j=1,\ldots,J$ and $\sigma^{2}$, respectively. Correspondingly we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. The following theorem presents the null limiting properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the settings for model (3.9).
Theorem 4.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the null hypothesis $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we still have
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}.$
Theorem 4 is parallel with Theorem 1 and Corollary 1 except for extra conditions regarding the group sample sizes and the number of groups. These additional conditions are imposed for technical considerations and to account for the heterogeneity across group means from model (3.9). We illustrate the intuition behind the conditions as follows. First, we require that the sample size of each group, $N_{j}$, tends to infinity. For model (3.9) where $\mu_{j}$ is allowed to be distinct, only the observations in group $j$ can be used to estimate $\mu_{j}$, thus $N_{j}\to\infty$ is necessary for the consistency of $\hat{\mu}_{j}$. Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must be satisfied jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$, which rules out the case that the number of groups grows faster than the group sample sizes. In contrast, for model (3.1) in the previous section, only the total sample size $N\to\infty$ suffices to derive the asymptotic properties regardless of the magnitude of $N_{j}$s and $J$. This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$. Thus, the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are all i.i.d., and we can just pool them together for efficient estimation and inference.
The following two theorems establish the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined by (3.3)) under the alternatives. The conditions are the same as those of Theorem 4, and the results are parallel with those in Theorems 2 and 3.
Theorem 5.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore,
$\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K},$
and
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right).$
Theorem 6.
Suppose Assumptions 3.1 and 3.2 hold. Then, under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right).$
Theorems 4, 5, and 6 provide an asymptotic $\chi^{2}$ test for the normality of model (3.9) and demonstrate its power properties. Such a test is quite similar to that in Section 3.1. When the group means in (3.9) are all equal in the sense that $\mu_{1}=\ldots=\mu_{J}=\mu$, the methodology in this section can be reduced to that in Section 3.1 for model (3.1), which validates the unified inferential framework and reflects the effects of group heterogeneity in means.
3.3 Same mean and different variances
Now we allow the group variances $\sigma$ in model (3.1) to be different among the groups, in the sense that
$Y_{ij}=\mu+\varepsilon_{ij}=\mu+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (3.10)
where $\{\sigma_{j}\}_{j=1}^{J}$ represents the potentially different standard deviations of each group. This time, we adopt the estimators $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}(Y_{ij}-\hat{\mu})^{2}$ for $\mu$ and $\sigma_{j}^{2},j=1,\ldots,J$, respectively. Note that the estimator of the population mean here is different from the sample mean in Section 3.1 due to the heterogeneity of variation across groups. Let $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu})/\hat{\sigma}_{j}$ and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. Define $p_{j}=\lim N_{j}/N$ and $q_{j}=\lim JN_{j}/N$ regarding the relative proportion among the group sample sizes. Here we further impose the following conditions for model (3.10): (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$. They both imply that the sample sizes and the standard deviations of each group are of the same magnitude.
Like the procedure in Sections 3.1 and 3.2 above, we start from the asymptotic behavior of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ here, see the following theorem.
Theorem 7.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under $H_{0}:e\sim\Phi(x)$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.11)
$+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
Note that the decomposition of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ in Theorem 7 is no longer the same as that in Theorems 1 and 4. The difference lies in the parameter estimation effect of $\hat{\mu}$ characterized by the term $c_{1k}N^{-1}(\sum_{\ell=1}^{J}p_{\ell}\sigma^{-1}_{\ell})\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\sigma_{j}e_{ij}/q_{j}$ on the right hand side of (3.11). Unlike the adoption of the sample mean for estimating $\mu$ in model (3.1), here we consider a new form of $\hat{\mu}$ for model (3.10), which leads to a different estimation effect term. Specially, when all groups share the same variance, i.e., $\sigma_{1}=\ldots=\sigma_{J}$, the term can be simplified as $c_{1k}N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}=-c_{1k}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}e_{ij}$; further if the group sample sizes are equal, i.e., $N_{1}=\ldots=N_{J}$, then the estimation effect reduces to that in Theorem 1, $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$. In contrast, the estimation effect of the variances is still aggregated as the term $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ regardless of the equality of group variances or group sample sizes.
The asymptotic normality of the vector $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$ can be derived from (3.11). Under the conditions of Theorem 7, by CLT,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{Z}_{j},$
where $\bm{Z}_{1},\ldots,\bm{Z}_{J}$ are independently distributed as
$\bm{Z}_{j}\sim\mathcal{N}_{K}\left(\bm{0},\bm{\Omega}_{K}^{(j)}\right),$
with $\bm{\Omega}_{K}^{(j)}=\left(\omega_{kl}^{(j)}\right)_{K\times K}$ given by
$\omega_{kl}^{(j)}$ $=\mathrm{E}\left\{\pi_{k}(Z)-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2k}}{2}\left(e^{2}-1\right)\right\}$
$\qquad\left\{\pi_{l}(Z)-c_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{c_{2l}}{2}\left(e^{2}-1\right)\right\}$
$=\delta_{kl}-\frac{2c_{1k}c_{1l}\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{c_{1k}c_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}-\frac{c_{2k}c_{2l}}{2}.$
Therefore, under $H_{0}$, we have
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right).$
By replacing $\sigma_{j}$, $p_{j}$ and $q_{j}$ by $\hat{\sigma}_{j}$, $\hat{p}_{j}=N_{j}/N$ and $\hat{q}_{j}=JN_{j}/N$ respectively, we obtain a consistent estimator of $\bm{\Omega}_{K}^{(j)}$, denoted as $\widehat{\bm{\Omega}}_{K}^{(j)}$. Correspondingly, the test statistic takes the form
$\hat{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}\hat{p}_{j}\widehat{\bm{\Omega}}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.12)
We also define the infeasible version,
$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right).$ (3.13)
Corollary 2.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under $H_{0}:e\sim\Phi(x)$,
[…]
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$.
Corollary 2 demonstrates that under the null hypothesis, both the feasible test statistic $\hat{\Psi}_{K}^{2}$ (in (3.12)) and its infeasible version $\tilde{\Psi}_{K}^{2}$ (in (3.13)) are asymptotically $\chi_{K}^{2}$ distributed after multiplied by $N$, the total sample size. Although the test statistics considered here take a slightly different form from those in the earlier setting, the limiting Chi-square distribution is retained and the convergence rate remains $N$, thereby reinforcing the validity of our unified theoretical framework. The following two theorems establish the asymptotic behavior of $\tilde{\Psi}_{K}^{2}$ under the fixed and local alternatives, which is also analogous to that of $\hat{\Psi}_{K}^{2}$ in Sections 3.1 and 3.2.
Theorem 8.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the alternative hypothesis $H_{1}$ in (3.4), for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (3.14)
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
$\tilde{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K},$
and
$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Upsilon}_{K}\bm{a}_{K}\right),$ (3.15)
where
$\bm{\Upsilon}_{K}=\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right)\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1},$
and $\bm{\Lambda}_{K}^{(j)}=\left(\lambda_{kl}^{(j)}\right)_{K\times K}$ is given by
$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
Theorem 9.
Suppose Assumptions 3.1 and 3.2 hold. Then, given the conditions of model (3.10), under the local alternative hypothesis $H_{1L}$ in (3.7), with $\delta_{N}=N^{-1/2}$, for $k=1,\ldots,K$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right])-c_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e_{ij}}{q_{j}}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\delta_{N}\Delta_{k}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$. Furthermore,
$N\tilde{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{\Delta}_{K}\right).$
Remark 2.
To conclude this section, we summarize the regularity conditions of the three models above in Table 1. On the one hand, one can find that more complex model structures lead to more restrictions on parameters and sample sizes. The heterogeneity across groups indeed poses additional challenges for theoretical analysis. On the other hand, under these mild conditions, an asymptotic Chi-square test with convergence rate $N$ can always be established for each of the three cases. This demonstrates that our proposed method possesses a unified theoretical framework while retaining the ability to capture the distinctive features of different ANOVA models.
Table 1: The regularity conditions for the three models
Model Conditions
Same mean & same variance (3.1) $N\to\infty$
Different means & same variance (3.9)
$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
$J=o(N^{1/2})$, $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$
Same mean & different variances (3.10)
$\min\{N_{1},\ldots,N_{J}\}\to\infty$,
$J=o(N^{1/2})$ , $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$,
$\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$,
$\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$
4 Data-driven choice of $K$
The methodology in Section 3 is based on the fact that the order of smooth alternatives (2.1), $K$, is a predetermined fixed integer before the tests are performed. In fact, the order $K$ has a key impact on the performance of smooth tests (Ledwina,, 1994; Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). If $K$ is large, the alternatives we test against contain redundant components that do not contribute to the test statistic markedly but increase the number of degrees of freedom. This will make smooth tests more omnibus and lead to power dilution. If $K$ is small, as described in Section 3, it is more likely to happen that $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$, which makes the tests exhibit no power. Therefore, it is crucial to select an appropriate $K$ before applying the test. In this section, we focus on the data-driven choice of $K$ for the three cases in Section 3.
In the literature of Neyman’s smooth tests, the spirit of data-driven methods was first introduced by Ledwina, (1994) for the problem of testing for uniformity. The data-driven procedure consists of two steps. First, Schwarz’s selection rule (also known as the Bayesian Information Criterion, BIC) is applied to find a suitable dimension of the smooth model for the data. Second, Neyman’s smooth test is performed within the selected model space, yielding a data-driven test statistic. Therefore, the data-driven smooth test serves as a kind of first selection followed by a more precise instrument, being Neyman’s test in the “right” direction. Desirable theoretical properties and extensive numerical results of the selection rule and the induced test statistic were shown in Kallenberg and Ledwina, 1995a and Kallenberg and Ledwina, 1995b , Inglot and Ledwina, (1996). Later, some modified versions of Schwarz’s selection rule were employed to address more complicated issues, such as testing for composite hypotheses (Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Inglot et al.,, 1997), testing for independence (Kallenberg and Ledwina,, 1999), two-sample testing problems (Janic-Wróblewska and Ledwina,, 2000), and so on. The modified selection rules utilize the smooth test statistics and circumvent the need to compute the maximized log-likelihood function of smooth models; thus, they are easier to implement.
Motivated by the strategies mentioned above, we adopt a modified Schwarz’s rule to determine the order $K$ for our problem. The selection rule is specified as
$\hat{K}=\min\left\{\mathop{\arg\max}\limits_{1\leq k\leq D}\left(N\hat{\Psi}_{k}^{2}-k\log N\right)\right\},$ (4.1)
where $D$ is some fixed positive integer, and $\hat{\Psi}_{k}^{2}$ takes the form of (3.3) or (3.12) up to specific scenarios. The corresponding data-driven test statistic is $N\Psi_{\hat{K}}^{2}$, where $\hat{K}$ is given by (4.1). The form of (4.1) is mainly inspired by Ducharme and Fontez, (2004), Ducharme and Lafaye de Micheaux, (2004), Kraus, (2007) and Duchesne et al., (2016) where the upper bound of selection $D$ is fixed. An alternative is to set $D=D(N)\to\infty$ as $N\to\infty$ (see, for example, Kallenberg and Ledwina, 1995a, ; Kallenberg and Ledwina, 1995b, ; Inglot and Ledwina,, 1996; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000). Although the diverging upper bound helps to improve the consistency of smooth tests theoretically, the rate of divergence is required to be quite slow, and the empirical power levels off rapidly as $D$ increases in their simulation studies. Therefore, our proposed selection rule (4.1) with fixed $D$ is reasonable and simple to implement in practice.
To establish the properties of $\hat{K}$ and $N\Psi_{\hat{K}}^{2}$, we need a revised version of the alternative hypotheses $H_{1}$ (3.4) and Assumption 3.2. Consider the following alternatives:
$H_{1}^{\prime}:\mathrm{E}[\pi_{1}(Z)]=\ldots=\mathrm{E}[\pi_{K_{0}-1}(Z)]=0,\mathrm{E}[\pi_{K_{0}}(Z)]\neq 0,\quad K_{0}\leq D.$ (4.2)
In other words, $H_{1}^{\prime}$ encompasses a broader range of alternatives than $H_{1}$ due to the replacement of the fixed order $K$ with the larger $D$. Analogously, the condition on the orthonormal system is now broadened to include all functions up to order $D$; see the revised assumption below.
Assumption 4.1.
For $k=1,\ldots,D$, $\pi_{k}(\cdot)$ are two times differentiable with derivatives $\dot{\pi}_{k}(\cdot)$ and $\ddot{\pi}_{k}(\cdot)$, and they are both bounded.
The following theorem demonstrates the properties of data-driven tests under the null and the revised alternatives.
Theorem 10.
Suppose Assumptions 3.1 and 4.1 hold.
  • (1).
    For model (3.1) in Section 3.1, as $N\to\infty$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
  • (2).
    For model (3.9) in Section 3.2, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$ and $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\hat{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
  • (3).
    For model (3.10) in Section 3.3, with additional conditions such that (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$, as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, under $H_{0}:e\sim\Phi(x)$, $\mathrm{P}(\hat{K}=1)\to 1$, $N\hat{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$ and $N\tilde{\Psi}_{\hat{K}}^{2}\to_{d}\chi_{1}^{2}$; under $H_{1}^{\prime}$ in (4.2), $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ and $\mathrm{P}(N\tilde{\Psi}^{2}_{\hat{K}}\leq x)\to 0$ for any $x\in\mathbb{R}$.
Theorem 10 presents the unified results of data-driven smooth tests for the three cases in Section 3. On the one hand, under the null hypothesis, the probability of $\{\hat{K}=1\}$ tends to $1$ asymptotically, which implies that the first-order smooth model fits $\{\hat{Z}_{ij}\}_{i=1,j=1}^{N_{j}.J}$ best among the $D$ candidate models. From the perspective of hypothesis testing, it is $\hat{\Psi}_{1}^{2}$ that is informative and adequate for diagnosing the uniformity/normality. On the other hand, the data-driven selection procedure helps identify the mechanism behind nonuniformity/nonnormality and leads to a more powerful test against a broader range of alternatives, $H_{1}^{\prime}$. These findings are parallel with existing results about data-driven smooth tests.
However, in practice, the $\chi_{1}^{2}$ limiting null distribution of data-driven smooth test statistics is often inaccurate in finite samples (Kallenberg and Ledwina, 1995b, ; Inglot et al.,, 1997; Kallenberg and Ledwina,, 1997; Kallenberg and Teresa,, 1997; Kallenberg and Ledwina,, 1999; Janic-Wróblewska and Ledwina,, 2000; Ducharme and Fontez,, 2004; Ducharme and Lafaye de Micheaux,, 2004; Kraus,, 2007; Duchesne et al.,, 2016). To overcome this issue in our problem, provided the nested orthonormal system (i.e. $\{\pi_{k}\}_{k=1}^{K}\subseteq\{\pi_{k}\}_{k=1}^{K+1}$ for $1\leq K\leq D-1$), we follow the instruction of Kallenberg and Ledwina, 1995b , Inglot et al., (1997), Kallenberg and Ledwina, (1997), Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007) and adopt a finite-sample approximation for the null distribution $N\hat{\Psi}_{\hat{K}}^{2}$:
[…]
Such an approximation primarily accounts for the selection uncertainty of $\hat{K}$. In finite samples, under the null, the empirical probability of observing $\{\hat{K}=1\}$ may not be exactly one, as the event $\{\hat{K}=2\}$ can still occur with a small but negligible probability. Accordingly, (4.3) is obtained from $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x)\approx\mathrm{P}(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1)+\mathrm{P}(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2)$. The reliability and effectiveness of $H(x)$ are demonstrated through simulation studies in the following section.
5 Simulation
In this section, we conduct Monte Carlo experiments to examine the finite-sample performance of the proposed tests developed in Sections 3 and 4. Following Neyman, (1937) and Song and Xiao, (2022), we adopt the orthonormal Legendre polynomials on the interval $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. Under this choice, the analytical properties of the constants $c_{1k}$ and $c_{2k}$ are provided in Proposition 1 of Duchesne et al., (2016), and numerical values for $1\leq k\leq 10$ are available in Table 1 of their Supplementary Material. From a computational perspective, the test statistic $N\hat{\Psi}_{K}^{2}$, introduced in Sections 3.1 and 3.2, can be computed using a simplified expression analogous to equation (14) in Duchesne et al., (2016), which leverages the closed-form of $\bm{\Sigma}_{K}^{-1}$ (see equations (12) and (13) therein). Since Section 3.1 is a special case of Section 3.2 or 3.3, we focus our evaluation on the settings described in Sections 3.2 and 3.3 through two numerical experiments.
In the first experiment, we generate observations with different means and the same variances as follows:
$Y_{ij}^{(0,1)}\sim\mathcal{N}\left(5j,4\right),\quad Y_{ij}^{(1,1)}\sim\chi_{2}^{2}+(5j-2),\quad i=1,\ldots,jm,\quad j=1,\ldots,5,$
where $\{Y_{ij}^{(0,1)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,1)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
$\mathrm{E}[Y_{ij}^{(0,1)}]=\mathrm{E}[Y_{ij}^{(1,1)}]=5j$ and $\operatorname{Var}[Y_{ij}^{(0,1)}]=\operatorname{Var}[Y_{ij}^{(1,1)}]=4$. The sample size parameter $m$ varies from 10 to 150 in increments of 10. We set the significance level $\alpha=5\%$ and perform $500$ replications. We evaluate the performance of the test statistic $N\hat{\Psi}_{K}^{2}$ introduced in Section 3.2, with $1\leq K\leq 5$, as well as the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ from Section 4, where $\hat{K}$ is determined by (4.1) with $D=5$. Table 2 reports the empirical rejection rates under $H_{0}$. For fixed values of $K=1,\ldots,5$, $N\hat{\Psi}_{K}^{2}$ maintains the nominal level well, even for small sample sizes, such as $m=10$ (corresponding to a total sample size of $N=150$). For the data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$, the results indicate that the limiting $\chi^{2}_{1}$ distribution does not adequately approximate the null distribution, whereas the revised distribution $H(x)$ achieves better size control. The empirical frequency of $\hat{K}$ under $H_{0}$ and $H_{1}$ is presented in Tables 3 and 4, respectively. Under $H_{0}$, the frequency of $\{\hat{K}=1\}$ approaches 1 as $m$ increases, while under $H_{1}$, $\hat{K}$ is consistently greater than 1, which aligns with our theoretical expectations. Additionally, Figure 1 displays the sample means and corresponding error bars (sample mean $\pm$ sample standard deviation) of $\hat{K}$ across different values of $m$. Here, under our data-generating processes, under $H_{0}$, the sample mean of $\hat{K}$ remains close to 1 regardless of the sample size, and the variability diminishes as $m$ increases, indicating consistent model selection under the null. Under $H_{1}$, $\hat{K}$ increases steadily and converges to the maximum selection range $D=5$ as $m$ grows, with reduced variability. These findings align with our theoretical results and demonstrate the adaptiveness of the data-driven procedure.
Table 2: Empirical rejection rates under $H_{0}$ in the first experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
$10$ 0.050 0.046 0.048 0.048 0.052 0.076 0.050
2020 0.072 0.060 0.050 0.046 0.050 0.054 0.048
$30$ 0.054 0.046 0.054 0.038 0.056 0.074 0.062
$40$ 0.046 0.048 0.060 0.054 0.048 0.054 0.050
$50$ 0.044 0.060 0.052 0.052 0.048 0.040 0.038
$60$ 0.066 0.044 0.048 0.052 0.058 0.048 0.042
$70$ 0.044 0.048 0.038 0.046 0.048 0.072 0.064
$80$ 0.050 0.044 0.048 0.038 0.042 0.074 0.068
$90$ 0.040 0.060 0.046 0.054 0.046 0.048 0.042
$100$ 0.062 0.056 0.042 0.050 0.048 0.058 0.052
$110$ 0.060 0.046 0.066 0.052 0.050 0.068 0.054
$120$ 0.034 0.044 0.038 0.038 0.036 0.058 0.056
$130$ 0.038 0.050 0.042 0.052 0.050 0.058 0.048
$140$ 0.038 0.050 0.048 0.042 0.044 0.044 0.040
$150$ 0.064 0.056 0.048 0.058 0.058 0.046 0.042
Table 3: Empirical frequency of $\hat{K}$ under $H_{0}$ in the first experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.976 0.024 0 0 0
2020 0.974 0.018 0.006 0 0.002
$30$ 0.986 0.014 0 0 0
$40$ 0.990 0.010 0 0 0
$50$ 0.998 0.002 0 0 0
$60$ 0.988 0.012 0 0 0
$70$ 0.992 0.008 0 0 0
$80$ 0.990 0.008 0.002 0 0
$90$ 0.988 0.012 0 0 0
$100$ 0.996 0.004 0 0 0
$110$ 0.994 0.004 0.002 0 0
$120$ 0.992 0.006 0.002 0 0
$130$ 0.998 0.002 0 0 0
$140$ 0.992 0.008 0 0 0
$150$ 0.996 0.004 0 0 0
Table 4: Empirical frequency of $\hat{K}$ under $H_{1}$ in the first experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.032 0.004 0 0.418 0.546
2020 0 0 0 0.314 0.686
$30$ 0 0 0 0.174 0.826
$40$ 0 0 0 0.098 0.902
$50$ 0 0 0 0.064 0.936
$60$ 0 0 0 0.036 0.964
$70$ 0 0 0 0.022 0.978
$80$ 0 0 0 0.018 0.982
$90$ 0 0 0 0.008 0.992
$100$ 0 0 0 0.008 0.992
$110$ 0 0 0 0.002 0.998
$120$ 0 0 0 0 1
$130$ 0 0 0 0 1
$140$ 0 0 0 0.002 0.998
$150$ 0 0 0 0 1
Figure 1: The sample means and error bars of $\hat{K}$ in the first experiment
Figure 2: The sample means and error bars of $\hat{K}$ in the second experiment
In the second experiment, we generate observations with the same means and different variances as follows:
$Y_{ij}^{(0,2)}\sim\mathcal{N}\left(8,j^{2}\right),\ Y_{ij}^{(1,2)}\sim\mathcal{U}\left[8-\sqrt{3}j,8+\sqrt{3}j\right],\ i=1,\ldots,jm,\ j=1,\ldots,5,$
where $\{Y_{ij}^{(0,2)}\}_{i=1,j=1}^{jm,5}$, $\{Y_{ij}^{(1,2)}\}_{i=1,j=1}^{jm,5}$ are generated under $H_{0}$ and $H_{1}$ respectively. Note that
$\mathrm{E}[Y_{ij}^{(0,2)}]=\mathrm{E}[Y_{ij}^{(1,2)}]=8$ and $\operatorname{Var}[Y_{ij}^{(0,2)}]=\operatorname{Var}[Y_{ij}^{(1,2)}]=j^{2}$. We adopt the same settings as in the first experiment. The results are analogous to those observed previously; see Tables 5, 6, and 7, as well as Figure 2. However, from Table 5, we observe that under $H_{1}$, the test statistic $N\hat{\Psi}_{1}^{2}$ fails to demonstrate power. This is because $\mathrm{E}[\pi_{1}(Z)]=\mathrm{E}[\pi_{1}(\Phi((Y_{ij}^{(1,2)}-8)/j))]=0$, indicating that $N\hat{\Psi}_{K}^{2}$ may lack consistency against certain directions of alternatives, particularly when $K$ is small. Correspondingly, both Table 6 and Figure 2 show that under $H_{1}$, the event $\{\hat{K}=1\}$ does not occur, which is consistent with Theorem 10 (3), where it is established that $\mathrm{P}(\hat{K}\geq K_{0})\to 1$ under $H^{\prime}_{1}$ with $K_{0}=2$.
To summarize the simulation results, we conclude that the proposed test statistic $N\hat{\Psi}_{K}^{2}$ maintains the nominal significance level well under $H_{0}$, and demonstrates high power under the alternatives unless $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$. The data-driven test statistic $N\hat{\Psi}_{\hat{K}}^{2}$ using the approximated null distribution $H(x)$ performs reliably in controlling the Type  error and achieves excellent power performance.
Table 5: Empirical rejection rates in the second experiment, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$).
$H_{0}$ $H_{1}$
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
$10$ 0.034 0.044 0.040 0.030 0.028 0.062 0.042 0.040 0.998 0.998 0.996 0.994 0.998 0.998
2020 0.062 0.064 0.060 0.056 0.058 0.060 0.050 0.034 1 1 1 1 1 1
$30$ 0.040 0.040 0.038 0.042 0.048 0.060 0.044 0.034 1 1 1 1 1 1
$40$ 0.038 0.042 0.044 0.028 0.032 0.052 0.040 0.024 1 1 1 1 1 1
$50$ 0.044 0.042 0.042 0.042 0.044 0.060 0.048 0.036 1 1 1 1 1 1
$60$ 0.062 0.080 0.052 0.064 0.074 0.082 0.066 0.016 1 1 1 1 1 1
$70$ 0.062 0.066 0.060 0.050 0.062 0.052 0.042 0.050 1 1 1 1 1 1
$80$ 0.046 0.048 0.042 0.058 0.036 0.040 0.038 0.042 1 1 1 1 1 1
$90$ 0.060 0.058 0.050 0.052 0.046 0.052 0.042 0.032 1 1 1 1 1 1
$100$ 0.052 0.038 0.046 0.058 0.042 0.054 0.048 0.018 1 1 1 1 1 1
$110$ 0.040 0.034 0.038 0.032 0.040 0.052 0.046 0.032 1 1 1 1 1 1
$120$ 0.056 0.044 0.066 0.046 0.046 0.066 0.058 0.040 1 1 1 1 1 1
$130$ 0.052 0.058 0.054 0.056 0.050 0.058 0.052 0.028 1 1 1 1 1 1
$140$ 0.050 0.060 0.054 0.042 0.040 0.074 0.068 0.028 1 1 1 1 1 1
$150$ 0.056 0.064 0.054 0.058 0.060 0.054 0.052 0.028 1 1 1 1 1 1
Table 6: Empirical frequency of $\hat{K}$ under $H_{0}$ in the second experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0.972 0.024 0.004 0 0
2020 0.986 0.012 0 0.002 0
$30$ 0.980 0.018 0.002 0 0
$40$ 0.992 0.008 0 0 0
$50$ 0.990 0.010 0 0 0
$60$ 0.996 0.004 0 0 0
$70$ 0.994 0.006 0 0 0
$80$ 0.996 0.004 0 0 0
$90$ 0.996 0.004 0 0 0
$100$ 0.990 0.010 0 0 0
$110$ 0.994 0.006 0 0 0
$120$ 0.992 0.006 0.002 0 0
$130$ 0.992 0.008 0 0 0
$140$ 0.986 0.014 0 0 0
$150$ 0.998 0.002 0 0 0
Table 7: Empirical frequency of $\hat{K}$ under $H_{1}$ in the second experiment
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
$10$ 0 0.990 0.008 0.002 0
2020 0 0.982 0.018 0 0
$30$ 0 0.990 0.008 0 0.002
$40$ 0 0.994 0.006 0 0
$50$ 0 0.996 0.004 0 0
$60$ 0 0.996 0.004 0 0
$70$ 0 0.994 0.006 0 0
$80$ 0 0.996 0.004 0 0
$90$ 0 1 0 0 0
$100$ 0 0.998 0.002 0 0
$110$ 0 0.996 0.004 0 0
$120$ 0 1 0 0 0
$130$ 0 1 0 0 0
$140$ 0 0.988 0.012 0 0
$150$ 0 1 0 0 0
6 Discussion of other ANOVA models
In this section, we briefly discuss smooth tests for normality of other ANOVA models beyond those in Section 3. The testing approach above is also applicable to these models.
6.1 One-way fixed effects model with different group means and variances
We first consider the one-way ANOVA model with different group means and different group variances, which is given by
$Y_{ij}=\mu_{j}+\varepsilon_{ij}=\mu_{j}+\sigma_{j}e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J.$ (6.1)
Here $\{\mu_{j}\}_{j=1}^{J},\{\sigma_{j}\}_{j=1}^{J}$ represent the potentially different group means and group standard deviations, respectively. Due to the heterogeneity across groups in model (6.1), the estimation of each group mean and group variance only depends on observations in the corresponding group, in the sense that $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij},\hat{\sigma}_{j}^{2}=N_{j}^{-1}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ for $j=1,\ldots,J$. Accordingly, we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}_{j}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}_{j}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. To implement the smooth test, one can follow the steps in Section 3 and derive a formula like regarding the estimation effects in $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij}),k=1,\ldots,K$. Then the test statistic is constructed as a quadratic form of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$, normalized by the inverse of its asymptotic covariance matrix. It is expected that an asymptotic $\chi^{2}$ test can be obtained under mild conditions for model (6.1), which leads to a unified testing framework for one-way fixed effects models.
6.2 One-way random effects model
Now we turn to the one-way random effects model. The observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are generated from following equation:
$Y_{ij}=\mu+A_{j}+\varepsilon_{ij}=\mu+A_{j}+\sigma e_{ij},\quad i=1,\ldots,N_{j},\quad j=1,\ldots,J,$ (6.2)
where the $\{A_{j}\}_{j=1}^{J}$ are i.i.d. random effects independent of the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, with $\mathrm{E}[A_{j}]=0$ and $\operatorname{Var}[A_{j}]=\sigma_{a}^{2}$. In this setting, we estimate $\sigma^{2}$ by
$\hat{\sigma}^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\bar{Y}_{\cdot j}\right)^{2}=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\hat{\varepsilon}_{ij}^{2},$
where $\bar{Y}_{\cdot j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\varepsilon}_{ij}=Y_{ij}-\bar{Y}_{\cdot j}$. Note that $\bar{Y}_{\cdot j}$ coincides with $\hat{\mu}_{j}$ in Section 3.2, leading to an identical expression for $\hat{\sigma}^{2}$ here as that in Section 3.2, too. Therefore, the methodology in Section 3.2 can be directly applied to test the normality of $e_{ij}$ in (6.2), since the random effects $A_{j}$ are effectively eliminated by within-group centering via $Y_{ij}-\bar{Y}_{\cdot j}$.
6.3 Two-way models
Next, we extend the discussion to two-way ANOVA models. Suppose the observations
$\{Y_{ij_{1}j_{2}}\}_{i=1,j_{1}=1,j_{2}=1}^{N_{j_{1}j_{2}},J_{1},J_{2}}$ are generated from the two-way fixed effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+\alpha_{j_{1}}+\beta_{j_{2}}+\gamma_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.3)
subject to the identifiability constraints
$\sum_{j_{1}=1}^{J_{1}}\alpha_{j_{1}}=\sum_{j_{2}=1}^{J_{2}}\beta_{j_{2}}=\sum_{j_{1}=1}^{J_{1}}\gamma_{j_{1}j_{2}}=\sum_{j_{2}=1}^{J_{2}}\gamma_{j_{1}j_{2}}=0.$
Inspired by the methodology developed in Section 3, we can construct analogous estimators for $\mu$, $\{\alpha_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{\beta_{j_{2}}\}_{j_{2}=2}^{J_{2}}$, $\{\gamma_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and $\sigma^{2}$, and derive the corresponding test statistic. The theoretical framework in Section 3 remains valid in this setting.
Similarly, for the two-way random effects model:
$Y_{ij_{1}j_{2}}=\mu_{j_{1}j_{2}}+\varepsilon_{ij_{1}j_{2}}=\left(\mu+A_{j_{1}}+B_{j_{2}}+D_{j_{1}j_{2}}\right)+\sigma e_{ij_{1}j_{2}},$ (6.4)
where the i.i.d. zero-mean random effects $\{A_{j_{1}}\}_{j_{1}=1}^{J_{1}}$, $\{B_{j_{2}}\}_{j_{2}=1}^{J_{2}}$, $\{D_{j_{1}j_{2}}\}_{j_{1}=1,j_{2}=1}^{J_{1},J_{2}}$, and the random errors $\{e_{ij}\}_{i=1,j=1}^{N_{j},J}$, are mutually independent. To test the normality of $e_{ij}$ in (6.4), one can mimic the strategy used in Section 6.2 to eliminate the random effects first through within-cell centering, and then construct the test statistic based on the resulting residuals.
To conclude, the methodology proposed in Section 3 can be seamlessly extended to more general ANOVA settings. This unified framework offers a flexible and effective approach to testing for normality in a wide range of linear models.
7 Concluding remarks
In this paper, we establish Neyman’s smooth tests for assessing the normality assumption in ANOVA models. For the proposed tests in three types of one-way fixed effects models, we demonstrate the asymptotic properties and validate the finite-sample performance through extensive simulations. Our approach provides a reliable tool for validating the normality assumption in diverse ANOVA settings.
Several directions remain open for future research. First, although we have developed tests within the ANOVA framework, the true data-generating mechanism may not always be known a priori. A promising approach is to conduct preliminary structural tests to identify the most suitable model before applying the proposed procedures. Second, the independence assumption on random errors can be relaxed to enable further theoretical analysis. For example, repeated clinical trials frequently exhibit autocorrelated errors (Ma et al.,, 2012; Schober and Vetter,, 2018; Langenberg et al.,, 2022). In such cases, our smooth test framework could potentially be adapted by incorporating estimated correlation structures. Finally, for the data-driven selection rule, it is intriguing to explore the case where the upper bound $D=D(N)\to\infty$ as $N\to\infty$ from the theoretical perspective.
Appendix A Appendix
We provide the proof of our main theoretical results and some practical issues in this Appendix. The symbol $\lesssim$” means that the left side is bounded by a positive constant times the right side. The symbol $\asymp$” means that both sides are asymptotically equivalent.
Proof of Theorem 1.
We first study the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$. Note that
$\hat{\mu}-\mu=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\mu\right)=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{\sigma}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.1)
Under $H_{0}$, $e_{ij}\sim\mathcal{N}(0,1)i.i.d$., then $\sqrt{N}(\hat{\mu}-\mu)\sim\mathcal{N}(0,\sigma^{2})$, and $\hat{\mu}-\mu=O_{p}(N^{-1/2})$ as $N\to\infty$. Besides,
$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\left(\hat{\mu}-\mu\right)^{2}$
$=$ $\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+O_{p}\left(\frac{1}{N}\right).$ (A.2)
Under $H_{0}$, by CLT, $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$ as $N\to\infty$, then $\hat{\sigma}^{2}-\sigma^{2}=O_{p}(N^{-1/2})$.
Now we deal with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$. By second order Taylor expansion of $\pi_{k}(\hat{Z}_{ij})$ with respect to $\hat{Z}_{ij}$ at $Z_{ij}$, we obtain
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left(\hat{Z}_{ij}-Z_{ij}\right)$
$+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2},$ (A.3)
[…]
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)+\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2},$ (A.4)
where $\tilde{\varepsilon}_{ij}/\tilde{\sigma}$ an intermediate point between $^{\hat{\varepsilon}_{ij}}/{\hat{\sigma}}$ and $\varepsilon_{ij}/\sigma$, whose exact value may vary across different scenarios.
We focus on the first term in (A.4). Straightforward calculation leads to
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
$=$ $-\frac{\hat{\mu}-\mu}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\frac{\hat{\mu}-\mu}{\sigma}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]-\frac{\hat{\sigma}^{2}-\sigma^{2}}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]$
$+o_{p}\left(\hat{\mu}-\mu\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right),$ (A.5)
where the last step follows due to the law of large numbers (LLN) for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)]+o_{p}(1)$ and $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]+o_{p}(1)$ as $N\to\infty$, as well as the consistency of $\hat{\sigma}$ to $\sigma$ from the arguments above. Under $H_{0}$, through integration by parts, we have
$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=\int_{0}^{1}\pi_{k}(z)\Phi^{-1}(z)\mathrm{d}z=c_{1k},$
and
$\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=\int_{0}^{1}\pi_{k}(z)\left[\left(\Phi^{-1}(z)\right)^{2}-1\right]\mathrm{d}z=\int_{0}^{1}\pi_{k}(z)\left(\Phi^{-1}(z)\right)^{2}\mathrm{d}z=c_{2k},$
where we use the fact that $\mathrm{d}\Phi^{-1}(z)=\mathrm{d}z/\phi(\Phi^{-1}(z))$ and $\int_{0}^{1}\pi_{k}(z)\mathrm{d}z=0$. Thus, along with (A.1) and (A.2), (A.5) can be re-expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})-c_{1k}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{c_{2k}}{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.6)
It suffices to show that
$R_{N1}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\ddot{\pi}_{k}\left(\tilde{Z}_{ij}\right)\left(\hat{Z}_{ij}-Z_{ij}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.7)
and
$R_{N2}\equiv\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\dot{\phi}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.8)
For $R_{N1}$, by the boundedness of $\ddot{\pi}_{k}$ (in Assumption 3.2) and $\phi$,
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left[\Phi\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)-\Phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\right]^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\phi^{2}\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$=\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}.$
By the inequality $(a-b)^{2}\leq 2(a^{2}+b^{2})$ for any $a,b\in\mathbb{R}$,
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\asymp$ $\left(\hat{\mu}-\mu\right)^{2}+(\hat{\sigma}^{2}-\sigma^{2})^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
$=$ $O_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{N}\right)\left(\sigma^{2}+o_{p}(1)\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
where the last step is the result of (A.1), (A.2) and LLN for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}=\sigma^{2}+o_{p}(1)$. For $R_{N2}$, we have
$|R_{N2}|$ $=\left|\frac{1}{2N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\left[-\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\phi\left(\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right)\right]\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}\right|$
$\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left|\frac{\tilde{\varepsilon}_{ij}}{\tilde{\sigma}}\right|\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{2}$
$\leq\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right)^{1/2}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}\right)^{1/2},$
where the first “==” is due to $\dot{\phi}(x)=-x\phi(x)$, and the last $\leq$” is due to Cauchy–Schwarz inequality. Similar arguments as above yield that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{{\varepsilon}_{ij}}{\sigma}\right)^{4}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
$\asymp$ $\left(\hat{\mu}-\mu\right)^{4}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)=O_{p}\left(\frac{1}{N^{2}}\right),$
and
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}$ $\leq\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{{\varepsilon}_{ij}}{{\sigma}}\right)^{2}$
$=1+\left(1+o_{p}(1)\right)=O_{p}(1).$ (A.9)
Thus, $|R_{N2}|=O_{p}(N^{-1})=o_{p}(N^{-1/2})$. Hence (A.7) and (A.8) are verified. Combining (A.3), (A.4), (A.5), (A.6), (A.7) and (A.8), we complete the proof of Theorem 1. $\square$
Proof of Theorem 2.
The proof of Theorem 2 is similar to that of Theorem 1. Under $H_{1}$, (A.1) still holds, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ since $e_{ij}\sim F\neq\Phi$. Analogously, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\sigma^{2}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$.
We start from (A.3). Under $H_{1}$, the arguments and results in (A.3), (A.4) and (A.5) still hold. Unlike (A.6), here in the context of $H_{1}$ the last line of (A.5) should be expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})-d_{1k}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{d_{2k}}{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.10)
where we replace $c_{1k}$ with $d_{1k}$ and $c_{2k}$ with $d_{2k}$. Besides, the remainder terms $R_{N1}$ in (A.7) and $R_{N2}$ in (A.8) are still $O_{p}(N^{-1/2})$ through similar derivation above. Combining (A.3), (A.4), (A.5), (A.7), (A.8) and (A.10), we obtain (3.5) in Theorem 2. From (3.5), we know that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1),$
as $N\to\infty$, which implies $\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$. Furthermore,
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\bm{\Sigma}_{K}^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]$
$-\sqrt{N}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$
$=$ $2a^{\top}\bm{\Sigma}_{K}^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}\bm{\Sigma}_{K}^{-1}$
$\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.11)
The following part demonstrates the asymptotic normality of
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right).$
Indeed, the $k$-th component in it is
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\pi_{k}(\hat{Z}_{ij})-a_{k}\right)=\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\zeta_{k}(e_{ij})-\mathrm{E}\left[\zeta_{k}(e_{ij})\right]\right)+o_{p}(1),$
where
$\zeta_{k}(e)\equiv\pi_{k}(\Phi(e))-d_{1k}e-\frac{d_{2k}}{2}(e^{2}-1).$
Then, by CLT,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Xi}_{K}\right),$
with $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ given by
$\xi_{kl}=$ $\mathrm{E}\left[\zeta_{k}(e)\zeta_{l}(e)\right]$
$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
Along with (A.11), (3.6) in Theorem 2 is validated. This completes the proof of Theorem 2. $\square$
Proof of Theorem 3.
The proof of Theorem 3 is similar to that of Theorems 1 and 2. We just follow the arguments in the previous proof and carefully specify the behavior of the estimation effects. Under $H_{1L}$ in (3.7), the PDF of $e$ is given by
$f(x)=(1-\delta_{N})\phi(x)+\delta_{N}q(x).$
We first examine the convergence of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$ as before. For $\hat{\mu}-\mu$, (A.1) still holds under $H_{1L}$, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ as $N\to\infty$. For $\hat{\sigma}^{2}-\sigma^{2}$, under $H_{1L}$, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$  [^1] Here under $H_{1L}$, the asymptotic normality of $\sqrt{N}(\hat{\mu}-\mu)$ and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})$ are both derived from CLT for triangular arrays since the distribution of $e_{ij}$ changes over $N$. It can be verified that the conditions of Lindeberg–Feller CLT hold for both random sequences, see, for example, Proposition 2.27 of Van der Vaart, (2000) . since
$\mathrm{E}\left[e^{4}\right]=$ $\int_{-\infty}^{\infty}x^{4}f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}x^{4}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}x^{4}q(x)\mathrm{d}x$
$=$ $3+o(1).$
Next, we clarify the relationship of the denoted constants. Recall the definitions of the constants $d_{1k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$ and $d_{2k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)e]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$. Then under $H_{1L}$, as $N\to\infty$, we have
$d_{1k}=$ $\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi^{2}\left(x\right)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)q(x)\mathrm{d}x$
$=$ $c_{1k}+o(1),$ (A.12)
$d_{2k}=$ $\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))x\phi^{2}\left(x\right)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))x\phi\left(x\right)q(x)\mathrm{d}x$
$=$ $c_{2k}+o(1),$ (A.13)
and
$\mathrm{E}[\pi_{k}(Z)]=$ $\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)f(x)\mathrm{d}x$
$=$ $\left(1-\delta_{N}\right)\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)q(x)\mathrm{d}x$
$=$ $\delta_{N}\Delta_{k}.$ (A.14)
Now we start from (A.3) again. Under $H_{1L}$, we still have (A.3), (A.4) and (A.5). (A.6) also holds as a result of (A.12) and (A.13). The remainder terms $R_{N1}$ in (A.7) and $R_{N2}$ in (A.8) are still $O_{p}(N^{-1/2})$. Combining (A.3), (A.4), (A.5), (A.6), (A.7), (A.8) and (A.14), we obtain (3.8) in Theorem 3.
It suffices to derive the asymptotic normality of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$. Under $H_{1L}$, straightforward calculation yields that
$\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]=$ $\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{l}(\Phi(x))f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{k}(\Phi(x))\phi(x)\mathrm{d}x$
$+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{k}(\Phi(x))q(x)\mathrm{d}x$
$=$ $\delta_{kl}+o(1),$
$d_{3k}=$ $\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xf(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xq(x)\mathrm{d}x$
$=$ $c_{1k}+o(1),$
$d_{4k}=$ $\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}q(x)\mathrm{d}x$
$=$ $c_{2k}+o(1),$
and
$\mathrm{E}\left[e^{3}\right]=$ $\int_{-\infty}^{\infty}x^{3}f(x)\mathrm{d}x$
$=$ $(1-\delta_{N})\int_{-\infty}^{\infty}x^{3}\phi(x)\mathrm{d}x+\delta_{N}\int_{-\infty}^{\infty}x^{3}q(x)\mathrm{d}x$
$=$ $o(1).$
Thus, by (3.8) and CLT,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
$\to_{d}\mathcal{N}_{K}\left(\bm{\Delta}_{K},\bm{\Sigma}_{K}\right),$
since the asymptotic covariance matrix is still $\bm{\Sigma}_{K}$ from the calculation above. Besides,
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
by definition of the non-central $\chi^{2}$ distribution. This completes the proof of Theorem 3. $\square$
Proof of Theorem 4, 5 and 6.
The proof is similar to that of Theorems 1, 2, and 3. Note that in the case of different means and same variance we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ without distinction of the null and alternatives. Our goal is to prove that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]e_{ij}-\frac{\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.15)
We first focus on the estimation effects of $\hat{\mu}_{j}-\mu_{j}$ and $\hat{\sigma}^{2}-\sigma^{2}$. For $\hat{\mu}_{j}-\mu_{j}$, by CLT, we have $\sqrt{N_{j}}(\hat{\mu}_{j}-\mu_{j})\to_{d}\mathcal{N}(0,\sigma^{2})$ for each $1\leq j\leq J$ as $\min\{N_{1},\ldots,N_{J}\}\to\infty$. Besides, since
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\right]=\frac{J}{N}\sigma^{2},$ (A.16)
we have $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}=O_{p}(J/N)=o_{p}(N^{-1/2})$ from $J=o(N^{1/2})$. Analogously, by
$\mathrm{E}\left[\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{1}{N_{j}^{4}}\left[3N_{j}(N_{j}-1)+N_{j}\mathrm{E}\left[e^{4}\right]\right]=\frac{3}{N_{j}^{2}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{3}},$
we obtain
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}\right]$
$=$ $\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\left(\frac{3}{N_{j}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{2}}\right),$ (A.17)
which implies $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}=N^{-1}O_{p}(\sum_{j=1}^{J}N_{j}^{-1})=o_{p}(N^{-1})$ since $\sum_{j=1}^{J}N_{j}^{-1}=o_{p}(1)$. For $\hat{\sigma}^{2}-\sigma^{2}$, we have
$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}$
$=$ $\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
with the last step following (A.16). Note that the main term in the last line is the same as (A.2).
Next, we will show that
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.18)
Indeed, since
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)e_{i_{2}j},$
and
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}=\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}e_{i_{2}j},$
then we have
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{ij}$
$+\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
$\equiv$ $A_{N1}+A_{N2}.$
For $A_{N1}$, we have
$\mathrm{E}[A_{N1}]=\frac{J}{N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right],$
and
$\operatorname{Var}[A_{N1}]=\operatorname{Var}\left[\left(\dot{\pi}_{k}(Z)\phi\left(e\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}},$
then $A_{N1}=o_{p}(N^{-1/2})$ since $J=o(N^{1/2})$ as $N\to\infty$. For $A_{N2}$, we have $\mathrm{E}[A_{N2}]=0$, and
$\operatorname{Var}[A_{N2}]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\operatorname{Var}\left[\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}^{2}}\mathrm{E}\left[\left(\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}\right)^{2}\right]$
$=$ $\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}(N_{j}-1)}{N_{j}^{2}}\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)$
$=$ $\left(\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]+\mathrm{E}^{2}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\right)\frac{1}{N^{2}}\left(J-\sum_{j=1}^{J}\frac{1}{N_{j}}\right),$
then $A_{N2}=o_{p}(N^{-1/2})$. Hence, (A.18) is verified.
Now we start from (A.3). In the case of different means and same variance, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(\frac{{\varepsilon}_{ij}}{\sigma}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-e_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}$
$=$ $-\frac{1}{\hat{\sigma}}\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$-\frac{\hat{\sigma}^{2}-\sigma^{2}}{\hat{\sigma}(\hat{\sigma}+\sigma)}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$
$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}-\frac{1}{2\sigma^{2}}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
$+o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}\right)+o_{p}\left(\hat{\sigma}^{2}-\sigma^{2}\right)$
$=$ $-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$-\frac{1}{2N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.19)
where we apply (A.2), (A.18), LLN for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)e_{ij}$, and CLT for $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$ as $N\to\infty$.
It suffices to show $R_{N1}$ (in (A.7)) and $R_{N2}$ (in (A.8)) are $o_{p}(N^{-1/2})$ in order to obtain (A.15). The technique is similar to that in the proof of Theorem 1. For $R_{N1}$, we have
$|R_{N1}|$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{2}$
$\lesssim\frac{1}{N\sigma^{2}}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}+\left(\hat{\sigma}^{2}-\sigma^{2}\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}$
$\asymp O_{p}\left(\frac{J}{N}\right)+O_{p}\left(\frac{1}{N}\right)\left(\sigma^{2}+o_{p}(1)\right)=O_{p}\left(\frac{J}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
by (A.16) and the assumption $J=o(N^{1/2})$. For $R_{N2}$, we first have
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}$
$\lesssim$ $\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}^{2}-\sigma^{2})}{\hat{\sigma}\sigma(\hat{\sigma}+\sigma)}\right)^{4}\right]$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}+(\hat{\sigma}^{2}-\sigma^{2})^{4}\left(\mathrm{E}\left[e^{4}\right]+o_{p}(1)\right)$
$\asymp$ $o_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{N^{2}}\right)=o_{p}\left(\frac{1}{N}\right),$
by (A.17). Along with $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}^{2}=O_{p}(1)$ from (A.9), we obtain
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}}-\frac{\varepsilon_{ij}}{\sigma}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
Thus, (A.15) is verified. The form of (A.15) can be further refined under $H_{0}$, $H_{1}$ and $H_{1L}$ respectively. For example, under $H_{0}$, we can write $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ in (A.15), and then obtain the results of Theorem 4. $\square$
Proof of Theorem 7.
The technique is still the same as the previous ones. Recall that in the context of same mean and different variances, we denote $p_{j}=N_{j}/N$, $q_{j}=JN_{j}/N$ for $1\leq j\leq J$, and we further assume that as $\min\{N_{1},\ldots,N_{J}\}\to\infty$ and $J=o(N^{1/2})$, (i) there exist $0<\underline{\sigma}\leq\overline{\sigma}<\infty$ such that $\underline{\sigma}<\inf_{1\leq j\leq J}\sigma_{j}\leq\sup_{1\leq j\leq J}\sigma_{j}<\overline{\sigma}$; (ii) there exist $0<\underline{q}\leq\overline{q}<\infty$ such that $\underline{q}<\inf_{1\leq j\leq J}q_{j}\leq\sup_{1\leq j\leq J}q_{j}<\overline{q}$.
We first tackle the estimation effects of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$. For $\hat{\mu}-\mu$, note that $\hat{\mu}=J^{-1}\sum_{j=1}^{J}N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ here. Then,
$\hat{\mu}-\mu=\frac{1}{J}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=\frac{1}{J}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}.$ (A.20)
From (A.20) we obtain $\mathrm{E}[\hat{\mu}-\mu]=0$, and $\operatorname{Var}[\hat{\mu}-\mu]=J^{-2}\sum_{j=1}^{J}N_{j}^{-1}\sigma_{j}^{2}\asymp J^{-2}\sum_{j=1}^{J}N_{j}^{-1}=N^{-1}J^{-1}\sum_{j=1}^{J}q_{j}^{-1}=O_{p}(N^{-1})$. Then $\hat{\mu}-\mu=O_{p}(N^{-1/2})$. For $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}$, we have
$\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}$ $=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}\right)^{2}-\sigma_{j}^{2}\right)=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma_{j}^{2}\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}$ (A.21)
$=\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}$
$=O_{p}\left(\frac{1}{\sqrt{N_{j}}}\right)+O_{p}\left(\frac{1}{N}\right)+O_{p}\left(\frac{1}{\sqrt{N_{j}N}}\right)=O_{p}\left(\frac{1}{\sqrt{N_{j}}}\right),$
by CLT for $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}e_{ij}\to_{d}\mathcal{N}(0,1)$ and $N_{j}^{-1/2}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)\to_{d}\mathcal{N}(0,\mathrm{E}[e^{4}]-1)$ as $N_{j}\to\infty$ and (A.20).
We next prove some preliminary results such that
$\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.22)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.23)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.24)
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}=O_{p}(1),$ (A.25)
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$. For (A.22), straightforward calculation leads to
$\mathrm{E}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}},$
and
$\operatorname{Var}\left[\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right]=\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\sum_{j=1}^{J}\frac{p^{2}_{j}}{N_{j}\sigma^{2}_{j}}$
$=$ $\operatorname{Var}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{N_{j}}{\sigma^{2}_{j}}=O\left(\frac{1}{N}\right).$
Thus, (A.22) is verified. The derivation of (A.23) is analogous to that of (A.18) if we first rewrite $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ on the right side as
$\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}(e_{i_{2}j}^{2}-1),$
and so is (A.24). (A.25) is obtained by
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\}\right]=\mathrm{E}[\dot{\pi}_{k}(Z)\phi(e)e]\frac{1}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}=O(1),$
and
$\operatorname{Var}\left[\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi(e_{ij})e_{ij}\right]=O\left(\frac{1}{N}\right).$
Now we start from (A.3). In the case of the same mean and different variances, (A.3) and (A.4) still hold. The difference lies in that the first term of (A.4) should be further expressed as
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(\frac{\varepsilon_{ij}}{\sigma_{j}}\right)\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}$
$-\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}.$
$\equiv$ $D_{N1}-D_{N2}.$
For $D_{N1}$, under $H_{0}$, we have
$D_{N1}$ $=-\frac{\hat{\mu}-\mu}{N}\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\left(1+o_{p}(1)\right)$
$=-\frac{1}{N}\left(\sum_{j=1}^{J}\frac{N}{JN_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}\right)\left(\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}N}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}\right)\left(1+o_{p}(1)\right)$
$=-\frac{1}{N}\left(\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\sigma_{j}e_{ij}}{q_{j}}\right)\left(c_{1k}\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}+O_{p}\left(\frac{1}{\sqrt{N}}\right)\right)\left(1+o_{p}(1)\right)$
$=-\frac{c_{1k}}{N}\left[\sum_{j=1}^{J}\frac{p_{j}}{\sigma_{j}}\right]\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\sigma_{j}e_{ij}}{q_{j}}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.26)
where the second last line is followed by (A.22), and $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]=c_{1k}$ under $H_{0}$. The last line is followed by $N^{-1}\sum_{j=1}^{J}q_{j}^{-1}\sum_{i=1}^{N_{j}}\varepsilon_{ij}=o_{p}(1)$ using LLN. For $D_{N2}$, under $H_{0}$, we have
$D_{N2}=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left\{\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}\right$.
$\left.-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{2}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}$
$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{1}{\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}$
$-2\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}}{N_{j}\hat{\sigma}_{j}(\hat{\sigma}_{j}+\sigma_{j})}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
$=$ $\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{N_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}\left(e_{i_{2}j}^{2}-1\right)\right\}\left(1+o_{p}(1)\right)$
$+\left(\hat{\mu}-\mu\right)^{2}\frac{1}{2N}\sum_{j=1}^{J}\frac{1}{\sigma_{j}^{2}}\left\{\sum_{i=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{ij}\right)\phi\left(e_{ij}\right)e_{ij}\right\}\left(1+o_{p}(1)\right)$
$-\left(\hat{\mu}-\mu\right)\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}\sigma_{j}}\left\{\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\dot{\pi}_{k}\left(Z_{i_{1}j}\right)\phi\left(e_{i_{1}j}\right)e_{i_{1}j}e_{i_{2}j}\right\}$
$=$ $\frac{c_{2k}}{2}\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+o_{p}\left(\frac{1}{\sqrt{N}}\right),$ (A.27)
following (A.20), (A.21), (A.23), (A.24) and (A.25), as well as $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]=c_{2k}$ under $H_{0}$.
It suffices to show $R_{N1}$ (in (A.7)) and $R_{N2}$ (in (A.8)) are $o_{p}(N^{-1/2})$ in this case. For $R_{N1}$, referring to the arguments in the proof of Theorem 4, here we have
$|R_{N1}|\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{2}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2})}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{2}\equiv R_{N1}^{A}+R_{N1}^{B}.$
For $R_{N1}^{A}$, we obtain
$R_{N1}^{A}=\frac{(\hat{\mu}-\mu)^{2}}{N}\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}\asymp\frac{1}{N}O_{p}\left(\frac{1}{N}\right)O_{p}(N)=O_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{\sqrt{N}}\right),$
by (A.20) and
$\sum_{j=1}^{J}\frac{N_{j}}{\hat{\sigma}_{j}^{2}}=\sum_{j=1}^{J}\frac{N_{j}}{\sigma_{j}^{2}}(1+o_{p}(1))=O_{p}(N).$
For $R_{N1}^{B}$, we have
$R_{N1}^{B}\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{2}\right\}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\left(\frac{\sigma_{j}^{2}}{N_{j}}\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)+\left(\hat{\mu}-\mu\right)^{2}-2\left(\hat{\mu}-\mu\right)\frac{\sigma_{j}}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\lesssim$ $\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}\left(e_{ij}^{2}-1\right)\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)+\left(\hat{\mu}-\mu\right)^{4}\left(\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$+4\left(\hat{\mu}-\mu\right)^{2}\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\left(\sum_{i=1}^{N_{j}}e_{ij}^{2}\right)$
$\equiv$ $R_{N1}^{B,1}+\left(\hat{\mu}-\mu\right)^{4}R_{N1}^{B,2}+4\left(\hat{\mu}-\mu\right)^{2}R_{N1}^{B,3}.$
For $R_{N1}^{B,1}$, we rewrite it as
$R_{N1}^{B,1}=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\sum_{i_{3}=1}^{N_{j}}\left(e_{i_{1}j}^{2}-1\right)\left(e_{i_{2}j}^{2}-1\right)e_{i_{3}j}^{2}.$
Note that $\mathrm{E}[(e_{i_{1}j}^{2}-1)(e_{i_{2}j}^{2}-1)e_{i_{3}j}^{2}]=\mathrm{E}[(e^{2}-1)^{2}e^{2}]\mathbbm{1}(i_{1}=i_{2}=i_{3})+\mathrm{E}[(e^{2}-1)^{2}]\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$. Then,
$\mathrm{E}\left[R_{N1}^{B,1}\right]$ $=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{6}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[(e^{2}-1)^{2}e^{2}\right]+N_{j}(N_{j}-1)\mathrm{E}\left[(e^{2}-1)^{2}\right]\right)$
$=O\left(\frac{J}{N}\right),$
which implies $R_{N1}^{B,1}=o_{p}(N^{-1/2})$. Similarly, for $R_{N1}^{B,3}$, by $\mathrm{E}[e_{i_{1}j}e_{i_{2}j}e_{i_{3}j}^{2}]=\mathrm{E}[e^{4}]\mathbbm{1}(i_{1}=i_{2})+\mathbbm{1}(i_{1}=i_{2}\neq i_{3})$, we have
$\mathrm{E}\left[R_{N1}^{B,3}\right]=\frac{1}{N}\sum_{j=1}^{J}\frac{\sigma_{j}^{4}}{N_{j}^{2}}\left(N_{j}\mathrm{E}\left[e^{4}\right]+N_{j}(N_{j}-1)\right)=O\left(\frac{J}{N}\right),$
which implies $R_{N1}^{B,3}=o_{p}(N^{-1/2})$. For $R_{N1}^{B,2}$, straightforward calculation leads to
$\mathrm{E}\left[R_{N1}^{B,2}\right]=\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}\sigma_{j}^{2}\sum_{i=1}^{N_{j}}e_{ij}^{2}\right]=\frac{1}{N}\sum_{j=1}^{J}N_{j}\sigma_{j}^{2}=O(1),$
then $R_{N1}^{B,2}=O_{p}(1)$. Along with (A.20), we obtain $R_{N1}^{B}=o_{p}(N^{-1/2})$. Hence $R_{N1}=o_{p}(N^{-1/2})$. On the other hand, for $R_{N2}$, we first note that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}$ $\lesssim\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}-\varepsilon_{ij}}{\hat{\sigma}_{j}}\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\varepsilon_{ij}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)}{\hat{\sigma}_{j}\sigma_{j}(\hat{\sigma}_{j}+\sigma_{j})}\right)^{4}$
$\asymp\left(\hat{\mu}-\mu\right)^{4}+\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
$\asymp O_{p}\left(\frac{1}{N^{2}}\right)+o_{p}\left(\frac{1}{N}\right)=o_{p}\left(\frac{1}{N}\right),$
since
$\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{4}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}$
$\leq$ $\left\{\max_{1\leq j\leq J}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\right\}\left\{\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right\}$
$=$ $o_{p}\left(\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\sigma}_{j}^{2}-\sigma_{j}^{2}\right)^{2}\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\varepsilon_{ij}^{4}\right\}\right)$
$=$ $o_{p}\left(\frac{1}{\sqrt{N}}\right),$
where we use $\hat{\sigma}^{2}_{j}-\sigma^{2}_{j}=o_{p}(1)$ from (A.21), and the last step can be obtained through similar arguments above. Besides, $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\tilde{\varepsilon}_{ij}^{2}/\tilde{\sigma}_{j}^{2}=O_{p}(1)$ can be derived from the procedure in (A.9) analogously. Then we have
$\left|R_{N2}\right|\lesssim\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\frac{\tilde{\varepsilon}_{ij}^{2}}{\tilde{\sigma}_{j}^{2}}\right]^{1/2}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\frac{\hat{\varepsilon}_{ij}}{\hat{\sigma}_{j}}-\frac{\varepsilon_{ij}}{\sigma_{j}}\right)^{4}\right]^{1/2}=o_{p}\left(\frac{1}{\sqrt{N}}\right).$
Therefore, (3.11) is validated. This completes the proof of Theorem 7. $\square$
Proof of Theorem 8 and 9.
The proof of Theorem 8 is analogous to that of Theorem 2. Under $H_{1}$, in order to obtain (3.14), we just repeat the procedure in the proof of Theorem 7, and replace $c_{1k}$ with $d_{1k}$ in (A.26), $c_{2k}$ with $d_{2k}$ in (A.27). From (3.5) we know that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1).$
Thus,
$\tilde{\Psi}_{K}^{2}=\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left(\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)\right)\to_{p}\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}.$
Besides,
$\sqrt{N}\left(\tilde{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left\{\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]\right$.
$\left.\qquad-\bm{a}_{K}^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\bm{a}_{K}\right\}$
$=$ $2a^{\top}\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}$
$\left(\sum_{j=1}^{J}p_{j}\bm{\Omega}_{K}^{(j)}\right)^{-1}\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.28)
Under $H_{1}$, by CLT, we have
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right)-a_{1},\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)-a_{K}\right]^{\top}$
$\to_{d}\sum_{j=1}^{J}\sqrt{p_{j}}\bm{W}_{j},$
where $\bm{W}_{1},\ldots,\bm{W}_{J}$ are independently distributed as
$\bm{W}_{j}\sim\mathcal{N}_{K}\left(0,\bm{\Lambda}_{K}^{(j)}\right),$
with $\bm{\Lambda}_{K}^{(j)}=\left(\lambda_{kl}^{(j)}\right)_{K\times K}$ given by
$\lambda_{kl}^{(j)}=$ $\mathrm{E}\left\{\left(\pi_{k}(Z)-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2k}}{2}\left(e^{2}-1\right)\right\}$
$\quad\left\{\left(\pi_{l}(Z)-\mathrm{E}\left[\pi_{l}(Z)\right]\right)-d_{1l}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)\frac{\sigma_{j}e}{q_{j}}-\frac{d_{2l}}{2}\left(e^{2}-1\right)\right\}$
$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\frac{(d_{1k}d_{3l}+d_{1l}d_{3k})\sigma_{j}}{q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}+\frac{d_{1k}d_{1l}\sigma_{j}^{2}}{q_{j}^{2}}\left(\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\right)^{2}$
$+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{(d_{1k}d_{2l}+d_{1l}d_{2k})\sigma_{j}}{2q_{j}}\sum_{\ell=1}^{J}\frac{p_{\ell}}{\sigma_{\ell}}\mathrm{E}\left[e^{3}\right]$
$+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
Therefore,
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\sum_{j=1}^{J}p_{j}\bm{\Lambda}_{K}^{(j)}\right),$
and the asymptotic distribution of $\tilde{\Psi}_{K}^{2}$ in (3.15) can be derived from (A.28) through similar arguments above. This completes the proof of Theorem 8.
The proof of Theorem 9 can be referred to that of Theorem 3 due to the similar arguments and techniques used. To obtain the results in Theorem 9, following the proof of Theorem 8, it suffices to show that under $H_{1L}$, $\mathrm{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, $\mathrm{E}[\pi_{k}(Z)\pi_{l}(Z)]=\delta_{kl}+o(1)$, $d_{1k}=c_{1k}+o(1)$, $d_{2k}=c_{2k}+o(1)$, $d_{3k}=c_{1k}+o(1)$, $d_{4k}=c_{2k}+o(1)$, $\mathrm{E}[e^{3}]=o(1)$, and $\mathrm{E}[e^{4}]=3+o(1)$, which can be found in the proof of Theorem 3. $\square$
Proof of Theorem 10.
We focus on the first case in Theorem 10, and the latter two cases can be verified analogously. We primarily follow well-established steps in the literature, for example, the proof of Theorem 2 in Ducharme and Lafaye de Micheaux, (2004), Lemma 1 in Kraus, (2007), and Propositions 6 and 7 in Duchesne et al., (2016). Note that $\{\hat{K}=k\}\subseteq\{N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\}$ for $k^{\prime}\neq k$, then we have
$\mathrm{P}\left(\hat{K}=k\right)\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{k^{\prime}}-k^{\prime}\log N\right),\quad k^{\prime}\neq k.$ (A.29)
Under $H_{0}$, in order to show $\mathrm{P}(\hat{K}=1)\to 1$ as $N\to\infty$, we first show the contrary such that $\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 0$. In fact, from (A.29) we have
$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)$
$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{1}-\log N\right)$
$\leq$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq-\log N\right)$
$=$ $\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right).$
From Corollary 1 we know that under $H_{0}$, $N\hat{\Psi}^{2}_{k}\to_{d}\chi^{2}_{k}$ for each $k=1,\ldots,D$. Thus for $k=2,\ldots,D$, $\mathrm{P}(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N)\to 0$ as $N\to\infty$. Therefore,
$\sum_{k=2}^{D}\mathrm{P}\left(\hat{K}=k\right)\leq\sum_{k=2}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\geq(k-1)\log N\right)\to 0,$
as $N\to\infty$, then $\mathrm{P}(\hat{K}=1)=1-\sum_{k=2}^{D}\mathrm{P}(\hat{K}=k)\to 1$, and $N\hat{\Psi}^{2}_{\hat{K}}\to_{d}\chi^{2}_{1}$ follows.
Under $H_{1}^{\prime}$, (A.29) yields that
$\mathrm{P}\left(\hat{K}=k\right)$ $\leq\mathrm{P}\left(N\hat{\Psi}^{2}_{k}-k\log N\geq N\hat{\Psi}^{2}_{K_{0}}-K_{0}\log N\right)$
$=\mathrm{P}\left(\hat{\Psi}^{2}_{k}-k\frac{\log N}{N}\geq\hat{\Psi}^{2}_{K_{0}}-K_{0}\frac{\log N}{N}\right).$
From Theorem 2 we know that as $N\to\infty$, $\hat{\Psi}^{2}_{k}\to_{p}0$ for $k=1,\ldots,K_{0}-1$, $\hat{\Psi}^{2}_{K_{0}}\to_{p}\bm{a}_{K_{0}}^{\top}\bm{\Sigma}_{K_{0}}^{-1}\bm{a}_{K_{0}}>0$ and $\log N/N\to 0$. Thus $\mathrm{P}(\hat{K}=k)\to 0$ for $k=1,\ldots,K_{0}-1$, and
$\lim_{N\to\infty}\mathrm{P}\left(\hat{K}\geq K_{0}\right)=1-\lim_{N\to\infty}\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)=1.$
Then, for any $x\in\mathbb{R}$, as $N\to\infty$,
$\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)$ $=\sum_{k=1}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x,\hat{K}=k\right)$
$\leq\sum_{k=1}^{K_{0}-1}\mathrm{P}\left(\hat{K}=k\right)+\sum_{k=K_{0}}^{D}\mathrm{P}\left(N\hat{\Psi}^{2}_{k}\leq x\right)\to 0,$
since $\hat{\Psi}^{2}_{k}\to_{p}\bm{a}_{k}^{\top}\bm{\Sigma}_{k}^{-1}\bm{a}_{k}>0$ for $k=K_{0},\ldots,D$. Therefore, for any $x\in\mathbb{R}$,
$\lim_{N\to\infty}\mathrm{P}\left(N\hat{\Psi}^{2}_{\hat{K}}\leq x\right)=0,$
which means the test is consistent against the alternatives given by (4.2). $\square$
Intuition behind the revised limiting null distribution+ (4.3).
We mainly follow the arguments in Kallenberg and Ledwina, (1999), Janic-Wróblewska and Ledwina, (2000), Kraus, (2007), and provide a heuristic derivation of the approximation $H(x)$ for the first case since the latter two cases can be derived analogously. Note that
$\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x\right)\\ =\mathrm{P}\left(N\hat{\Psi}_{1}^{2}\leq x,\hat{K}=1\right)+\mathrm{P}\left(N\hat{\Psi}_{2}^{2}\leq x,\hat{K}=2\right)+\mathrm{P}\left(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3\right).$ (A.30)
The third term on the right-hand side of (A.30), $\mathrm{P}(N\hat{\Psi}_{\hat{K}}^{2}\leq x,\hat{K}\geq 3)$, can be neglected under $H_{0}$ (Kallenberg and Ledwina, 1995b, ). The event $\{\hat{K}=1\}$ is approximated by $\{N\hat{\Psi}^{2}_{1}-\log N\geq N\hat{\Psi}^{2}_{2}-2\log N\}=\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})\leq\log N\}$, and $\{\hat{K}=2\}$ is approximated by $\{N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})>\log N\}$.
We first investigate the asymptotic distribution of $(N\hat{\Psi}^{2}_{1},N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1}))^{\top}$. The limiting distributions of $N\hat{\Psi}^{2}_{1},N\hat{\Psi}^{2}_{2}$ are functions of a bivariate normal vector $(R_{1},R_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\bm{\Sigma}_{2})$. We generally denote elements of $\bm{\Sigma}_{2}$ as $\begin{pmatrix}a&b\\ b&c\end{pmatrix}$ and $\rho=b/\sqrt{ac}$. The distribution of $(R_{1},R_{2})^{\top}$ can be constructed through two independent standard normal variables $G_{1},G_{2}$ if we denote $\tilde{R}_{1}=\sqrt{a}(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})$ and $\tilde{R}_{2}=\sqrt{c}G_{2}$, then $(\tilde{R}_{1},\tilde{R}_{2})^{\top}\overset{d}{=}(R_{1},R_{2})^{\top}$. Thus, $N\hat{\Psi}^{2}_{1}\to_{d}R_{1}^{2}/a\overset{d}{=}\tilde{R}_{1}^{2}/a=(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2})^{2}$. For $N\hat{\Psi}_{2}^{2}$, straightforward computations yield that
$N\hat{\Psi}_{2}^{2}$ $\to_{d}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\bm{\Sigma}_{2}^{-1}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(\tilde{R}_{1},\tilde{R}_{2}\right)\begin{pmatrix}c&-b\\ -b&a\end{pmatrix}\left(\tilde{R}_{1},\tilde{R}_{2}\right)^{\top}$
$=\frac{1}{ac-b^{2}}\left(c\tilde{R}_{1}^{2}-2b\tilde{R}_{1}\tilde{R}_{2}+a\tilde{R}_{2}^{2}\right)$
$=G_{1}^{2}+G_{2}^{2}.$
Then $N(\hat{\Psi}_{2}^{2}-\hat{\Psi}_{1}^{2})\to_{d}(\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{2}$. Since $(\sqrt{1-\rho^{2}}G_{1}+\rho G_{2},\rho G_{1}-\sqrt{1-\rho^{2}}G_{2})^{\top}\sim\mathcal{N}_{2}(\bm{0},\mathbf{I}_{2})$, we obtain that $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ are both asymptotically $\chi_{1}^{2}$ distributed and are asymptotically independent.
Now we calculate the approximation $H(x)$. We will treat $H(x)$ separately for $x\leq\log N$, $\log N<x<2\log N$ and $x\geq 2\log N$. For $x\leq\log N$, note that
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)\approx\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}>\log N\right)=0,$
by $N\hat{\Psi}^{2}_{1}\geq 0$. Thus, we have the approximation
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)\approx\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]\equiv H(x),\quad x\leq\log N,$
from the asymptotic $\chi_{1}^{2}$ distribution and asymptotic independence of $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$. For $x\geq 2\log N$, we have
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)$
$\approx$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$=$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)-\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right).$
As $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ is approximately $\chi_{1}^{2}$ distributed, we obtain
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)\approx 2\left(1-\Phi\left(\sqrt{\log N}\right)\right)\approx 2\frac{\phi\left(\sqrt{\log N}\right)}{\sqrt{\log N}}=\sqrt{\frac{2}{\pi N\log N}},$
by the fact that $t\phi(t)/(t+1)\leq 1-\Phi(t)\leq\phi(t)/t$ for $t\to\infty$. Besides,
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x\right)$
$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>2\log N\right)$
$\approx$ $\exp\left\{-\frac{1}{2}\times 2\log N\right\}$
$=$ $\frac{1}{N},$
where we use the distribution function of the limiting $\chi^{2}_{2}$ for $N\hat{\Psi}^{2}_{2}$. Hence $\mathrm{P}(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$ converges to 0 much slower than $\mathrm{P}(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$, and thus the latter can be neglected. Therefore, for $x\geq 2\log N$, we have the approximation
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)+\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$\approx$ $\left[2\Phi\left(\sqrt{x}\right)-1\right]\left[2\Phi\left(\sqrt{\log N}\right)-1\right]+2\left[1-\Phi\left(\sqrt{\log N}\right)\right]$
$\equiv$ $H(x),\quad x\geq 2\log N.$
For $\log N<x<2\log N$, Kallenberg and Ledwina, 1995b suggested the linearization as follows
$H(x)=H(\log N)+\frac{x-\log N}{\log N}[H(2\log N)-H(\log N)],\quad\log N<x<2\log N.$
This completes the derivation of the approximation (4.3) for the null distribution of the test statistic $N\hat{\Psi}^{2}_{\hat{K}}$. $\square$
Practical issues in simulation studies.
In our simulation study, we adopt the orthonormal Legendre polynomials on $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. The properties of orthonormal Legendre polynomials lead to the following expressions for constants $c_{1k}$ and $c_{2k}$:
$c_{1k}$ $=\frac{\sqrt{2k+1}}{2^{k}}\,\sum_{j=0}^{[k/2]}(-1)^{j}\binom{k}{j}\binom{2(k-j)}{k}\int_{0}^{1}(2z-1)^{k-2j}\,\Phi^{-1}(z)\,\mathrm{d}z$
$=\sqrt{2k+1}\,\sum_{j=0}^{[k/2]}\left(\frac{-1}{4}\right)^{j}\binom{k}{j}\binom{2(k-j)}{k}\,\int_{-1/2}^{1/2}x^{k-2j}\,\Phi^{-1}\left(x+\frac{1}{2}\right)\,\mathrm{d}z,$
and
$c_{2k}=\sqrt{2k+1}\,\sum_{j=0}^{[k/2]}\left(\frac{-1}{4}\right)^{j}\binom{k}{j}\binom{2(k-j)}{k}\,\int_{-1/2}^{1/2}x^{k-2j}\,\left[\Phi^{-1}\left(x+\frac{1}{2}\right)\right]^{2}\,\mathrm{d}z.$
Note that $c_{2k}=0$ when $k$ is odd and $c_{1k}=0$ when $k$ is even. Therefore, in practice, we only compute the non-zero elements based on the above two formulae for $k=1,\ldots,K$ with $K\leq 5$.